OpenStudy (anonymous):

Which of the following equations is the result of completing the square on x^2 - 6x - 9 = 0?

OpenStudy (anonymous):

\[(x-3)^2=0\]

Parth (parthkohli):

That's right!

Parth (parthkohli):

Do they need the solution, or is that it? P.S. You were right there.

Parth (parthkohli):

Hey wait.

Parth (parthkohli):

\[(x - 3)^2 = x^2 - 6x + 9 \]

Parth (parthkohli):

So, there's a different way.

Parth (parthkohli):

\[x^2 - 6x - 9 = 0 \implies x^2 - 6x + 9 = 18 \implies (x - 3)^2 = 18\]\[\implies (x - 3)^2 - 18 = 0 \]